# QCD Phase Diagram and Astrophysical Implications

### v1

@@ 1 Important lessons we have learned from the heavy-ion collision
Although it is boring in phenomenology, the crossover is the most difficult to understand from the microscopic mechanism. In the duality region of the temperature, two descriptions in terms of hadrons and partons marginally work, and in turn, they nearly break down. Thus, besides the numerical simulations, there is no reliable theoretical device to investigate the matter properties in this intermediate region. Nevertheless, if one is interested in the thermodynamics only without caring about microscopic ingredients, one can construct it by a smooth interpolation between the HRG and the pQCD branches, as depicted in Fig. 1.
@@ 2 Not the best but the better among worse at high baryon density
Now, let us recall the lessons from the heavy-ion collision. In the past, many intriguing possibilities were considered, but after all, the most boring scenario turned out to be the case. Based on the wisdom we earned from high-temperature QCD matter, I would say, it should not be too conservative even if I postulate the least structured phase diagram, namely, only the crossover from hadronic matter to quark matter at high densities; see Fig. 2 for scenarios of the crossover and the first-order phase transition on the phase diagram. Because the duality plays such an essential role for deconfinement at high temperature, we must ask to ourselves; why not at high baryon density? Such an assumption of the least structured phase diagram may sound like the well-known principle of Occam’s razor, but I must say, the duality of deconfinement is a concrete physical concept and the crossover should conceal rich physics behind it.
@@ 3 Likely EoSs supporting the crossover but the first-order phase transitions popping up
I am not so negative against the scenario of the first-order phase transition. I am simply saying that the crossover scenario is a good starting point to draw the first approximation for the equation of state (EoS), and then one can engrave this base with more structures. To make my point clear, let us take a close look at the inferred EoSs from the machine-learning analysis. We have invested resources to establish a method to solve the inverse problem from the neutron star data on the radius-mass ($R$-$M$) plane to the EoS candidate. The technique was first developed by means of the Bayesian analysis. It is extremely nontrivial in this method of the Bayesian analysis how to evaluate the conditioned probability to find the observation data for the given EoS, that is, the likelihood. Once the EoS is specified, one line of the combination of $(R,M)$ of the neutron star is fixed. Then, suppose that this $(R,M)$ line is the true answer, how can we quantify the probability to obtain the observation data with error bars? Actually, the distribution of the observation data may be significantly different from the Gaussian, and there is no simple machinery to quantify the conditioned probability. Therefore, some sort of the Monte-Carlo simulation is indispensable, and the entire procedures are time consuming.
@@ 4 Trace anomaly approaching zero means what?
Figure 5: Conformality measure from the neutron star data (solid line) and the finite-temperature lattice-QCD data (dashed line). The horizontal axis is the energy in the unit of $\varepsilon_{0}=0.15\,\mathrm{GeV}/\mathrm{fm}^{3}$.
@@ 5 Scanning the phase transitions with the total mass of the binary neutron star system
It is the most ambitious attempt to utilize the gravitational wave signals for the purpose of the detection of the QCD phase transitions. For the moment, nobody knows the exact phase structure at high density, and my belief is the crossover from hadronic to quark matter as emphasized previously. As a matter of fact, the order of the phase transition itself is not so important in this context. If the pQCD estimate of the EoS is extrapolated toward the neutron star density, the pQCD-based EoS is very soft, and the speed of sound squared cannot be far from the conformal value. For the fate of the binary neutron star merger, it is essential that the EoS of dense matter after some phase transition or crossover is expected to be soft, and it is almost irrelevant how to transmute the matter with or without a discontinuous jump. The bottom line is the following: the QCD phase transition, which may be a first-order phase transition (Most et al., 2019; Weih et al., 2020) or a continuous crossover (Fujimoto et al., 2023), softens the EoS and the gravitational collapse to the blackhole is triggered by the onset of quark matter. The idea is schematically illustrated in Fig. 6. If the full shapes of the gravitational waves from the post-merger stage will be probed in the future, they would provide us with valuable information about the nature of the phase transitional change to quark matter. However, the nature is not so kind. After the merger, the system size is far smaller than the inspiral stage, and the amplitude of the gravitational waves is much reduced. The typical frequency range of the gravitational waves of our interest to investigate the QCD phase transition is around $\gtrsim 3\,\mathrm{kHz}$ and I am not very optimistic about the possibility to achieve the sensitivity of the detector in this frequency range. To confirm the QCD phase transition using the post-merger gravitational wave signals, we should be very lucky or we should complete the mission impossible to make a reliable and precise prediction or we should wait for decades till the detector sensitivity after several generations will be sufficiently improved.
@@
One workable breakthrough to overcome the sensitivity problem at high frequency band is the multi-messenger analysis. In the case of GW170817, an electromagnetic counterpart, AT2017gfo, identified as the kilonova has been observed. The extensive analysis of the kilonova gave us a lot of information about the r-process elements (Domoto et al., 2021). Here, I want to emphasize that the presence of the kilonova already imposes a condition onto the EoS candidate. The point is that the life time of the remnant after the merger should be long enough to acquire the disk mass in a way consistent with the estimated kilonova mass. Any phase transition scenario that cannot explain the kilonova mass must be strictly ruled out. If the luminosity of many observed kilonovae will be systematically investigated in the future, the phase transition scenarios can be discriminated even without looking at the post-merger signals. In fact, Fig. 7 shows the life time of the remnant after the merger as a function of the total mass of the binary neutron star system (Fujimoto et al., 2024a). CO and PT are two representative EoS scenarios. It is evident that the life time strongly depends on the EoS scenarios in a specific window of the total mass (labeled as “sweet spot” in Fig. 7). For example, according to Fig. 7, if a bright kilonova will be discovered for the merger with the total mass at $2.9M_{\odot}$, then the CO scenario must be denied. We note that, if the binary system is too light, the hypermassive neutron star lives long enough and the EoS differences are irrelevant. Also, if the binary system is too heavy, the remnant is immediately collapsed into the blackhole regardless of the EoS differences. The sweet spot is centered around twice of the typical neutron star mass $\sim 2.8M_{\odot}$, which was indeed the case for GW170817. In this way, the location and the strength of the phase transition including the crossover will be able to be constrained by the future observation in the total mass window of the sweet spot. In some events, the kilonova might be simply overlooked experimentally, but not always so. I am very optimistic about the systematic survey of the kilonova which will be able to render a verdict of the quark matter transition.

### v2

@@ 1 Important lessons we have learned from the heavy-ion collision
Although it is boring in phenomenology, the crossover as realized by the duality is the most difficult to understand from the microscopic mechanism. This makes a contrast to a theoretical description of quark hadron mixture; see, e.g., Blaschke et al. (2024) in a Beth-Uhlenbeck approach. In contrast, in the duality region of the temperature, two descriptions in terms of hadrons and partons marginally work, and in turn, they nearly break down. Thus, besides the numerical simulations, there is no reliable modeling of the duality to investigate the matter properties in this intermediate region (however, see Fujimoto et al. (2024c) for a recent development). Nevertheless, if one is interested in the thermodynamics only without caring about microscopic ingredients, one can construct it by a smooth interpolation between the HRG and the pQCD branches, as depicted in Fig. 1. See also Albright et al. (2014) for a hybrid matching prescription.
@@ 2 Not the best but the better among worse at high baryon density
Now, let us recall the lessons from the heavy-ion collision. In the past, many intriguing possibilities were considered, but after all, the most boring scenario turned out to be the case. Based on the wisdom we earned from high-temperature QCD matter, I would say, it should not be too conservative even if I postulate the least structured phase diagram, namely, only the crossover from hadronic matter to quark matter at high densities; see Fig. 2 for scenarios of the crossover and the first-order phase transition on the phase diagram. Because the duality plays such an essential role for deconfinement at high temperature, we must ask to ourselves; why not at high baryon density? Such an assumption of the least structured phase diagram may sound like the well-known principle of Occam’s razor, but I must say, the duality of deconfinement is a concrete physical concept and the crossover should conceal rich physics behind it. For possible theoretical justification, let me refer to the quark-hadron continuity scenario between a superfluid strange nuclear matter and a three-flavor color-superconducting matter (Schäfer and Wilczek, 1999; Alford et al., 1999; Hatsuda et al., 2006), which is extendable to the two-flavor case (Fujimoto et al., 2020). However, recent discussions of placing a superfluid vortex as a probe have revealed non-trivial topological nature of quark matter (Chatterjee et al., 2019; Alford et al., 2019; Cherman et al., 2019).
@@ 3 Likely EoSs supporting the crossover but the first-order phase transitions popping up
I am not so negative against the scenario of the first-order phase transition. I am simply saying that the crossover scenario is a good starting point to draw the first approximation for the equation of state (EoS), and then one can engrave this base with more structures. To make my point clear, let us take a close look at the inferred EoSs from the machine-learning analysis. We have invested resources to establish a method to solve the inverse problem from the neutron star data on the radius-mass ($R$-$M$) plane to the EoS candidate. The technique was first developed by means of the Bayesian analysis (Steiner et al., 2010; Ozel et al., 2016; Annala et al., 2020; Brandes et al., 2023). It is extremely nontrivial in this method of the Bayesian analysis how to evaluate the conditioned probability to find the observation data for the given EoS, that is, the likelihood. Once the EoS is specified, one line of the combination of $(R,M)$ of the neutron star is fixed. Then, suppose that this $(R,M)$ line is the true answer, how can we quantify the probability to obtain the observation data with error bars? Actually, the distribution of the observation data may be significantly different from the Gaussian, and there is no simple machinery to quantify the conditioned probability. Therefore, some sort of the Monte-Carlo simulation is indispensable, and the entire procedures are time consuming.
@@ 4 Trace anomaly approaching zero means what?
Figure 5: Conformality measure from the neutron star data (solid line) and the finite-temperature lattice-QCD data (dashed line). The horizontal axis is the energy density in the unit of $\varepsilon_{0}=0.15\,\mathrm{GeV}/\mathrm{fm}^{3}$.
@@ 5 Scanning the phase transitions with the total mass of the binary neutron star system
It is the most ambitious attempt to utilize the gravitational wave signals for the purpose of the detection of the QCD phase transitions. For the moment, nobody knows the exact phase structure at high density, and my belief is the crossover from hadronic to quark matter as emphasized previously. As a matter of fact, the order of the phase transition itself is not so important in this context. If the pQCD estimate of the EoS is extrapolated toward the neutron star density, the pQCD-based EoS is very soft, and the speed of sound squared cannot be far from the conformal value. For the fate of the binary neutron star merger, it is essential that the EoS of dense matter after some phase transition or crossover is expected to be soft, and it is almost irrelevant how to transmute the matter with or without a discontinuous jump. The bottom line is the following: the QCD phase transition, which may be a first-order phase transition (Most et al., 2019; Bauswein et al., 2019; Weih et al., 2020) or a continuous crossover (Fujimoto et al., 2023), softens the EoS and the gravitational collapse to the blackhole is triggered by the onset of quark matter. The idea is schematically illustrated in Fig. 6. If the full shapes of the gravitational waves from the post-merger stage will be probed in the future, they would provide us with valuable information about the nature of the phase transitional change to quark matter (for example, the shift of the peak frequency as discussed in Bauswein et al. (2019)). However, the nature is not so kind. After the merger, the system size is far smaller than the inspiral stage, and the amplitude of the gravitational waves is much reduced. The typical frequency range of the gravitational waves of our interest to investigate the QCD phase transition is around $\gtrsim 3\,\mathrm{kHz}$ and I am not very optimistic about the possibility to achieve the sensitivity of the detector in this frequency range. To confirm the QCD phase transition using the post-merger gravitational wave signals, we should be very lucky or we should complete the mission impossible to make a reliable and precise prediction or we should wait for decades till the detector sensitivity after several generations will be sufficiently improved.
@@
One workable breakthrough to overcome the sensitivity problem at high frequency band is the multi-messenger analysis. In the case of GW170817, an electromagnetic counterpart, AT2017gfo, identified as the kilonova has been observed. The extensive analysis of the kilonova gave us a lot of information about the r-process elements (Domoto et al., 2021). Here, I want to emphasize that the presence of the kilonova already imposes a condition onto the EoS candidate. The point is that the life time of the remnant after the merger should be long enough to acquire the disk mass in a way consistent with the estimated kilonova mass. Any phase transition scenario that cannot explain the kilonova mass must be strictly ruled out. If the luminosity of many observed kilonovae will be systematically investigated in the future, the phase transition scenarios can be discriminated even without looking at the post-merger signals. In fact, Fig. 7 shows the life time of the remnant after the merger as a function of the total mass of the binary neutron star system (Fujimoto et al., 2024a); see also the related discussions on the threshold mass in (Bauswein et al., 2020). CO and PT are two representative EoS scenarios. It is evident that the life time strongly depends on the EoS scenarios in a specific window of the total mass (labeled as “sweet spot” in Fig. 7). For example, according to Fig. 7, if a bright kilonova will be discovered for the merger with the total mass at $2.9M_{\odot}$, then the CO scenario must be denied. We note that, if the binary system is too light, the hypermassive neutron star lives long enough and the EoS differences are irrelevant. Also, if the binary system is too heavy, the remnant is immediately collapsed into the blackhole regardless of the EoS differences. The sweet spot is centered around twice of the typical neutron star mass $\sim 2.8M_{\odot}$, which was indeed the case for GW170817. In this way, the location and the strength of the phase transition including the crossover will be able to be constrained by the future observation in the total mass window of the sweet spot. In some events, the kilonova might be simply overlooked experimentally, but not always so. I am very optimistic about the systematic survey of the kilonova which will be able to render a verdict of the quark matter transition.